OpenStudy (lgbasallote):

How many 6 letter words can you create from the word ELEMENTARY? Note: The words do not have to be legitimate

OpenStudy (anonymous):

INFINATE!

OpenStudy (lgbasallote):

no...not really...there are finite letters

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

36

OpenStudy (anonymous):

What haha

OpenStudy (lgbasallote):

words

OpenStudy (lgbasallote):

yes exactly

OpenStudy (lgbasallote):

too bad you took it back though

OpenStudy (anonymous):

7C6 * 6! + 7C5 *6! + 7C4 * 6!/2! + 7C3 *6!/3!

OpenStudy (lgbasallote):

it's a pretty big number

OpenStudy (lgbasallote):

sadly...no @sauravshakya

OpenStudy (zarkon):

do you know the answer?

OpenStudy (lgbasallote):

yes

OpenStudy (zarkon):

8015

OpenStudy (lgbasallote):

i don't think so...how did you get it? i may be wrong

OpenStudy (zarkon):

\[_7P_6+ _7P_5+\frac{_7P_4}{2!}+\frac{_7P_3}{3!}\]

OpenStudy (zarkon):

I believe I am wrong now....give me a sec

OpenStudy (zarkon):

34650

OpenStudy (anonymous):

Can I ask a question?

OpenStudy (lgbasallote):

it's still too small...let me check

OpenStudy (zarkon):

I should prob write this down instead of just typing on my calc :)

OpenStudy (lgbasallote):

hmm actually that was too big

OpenStudy (anonymous):

First tell me how many 6 letter words can u create from a,b,c,d,e,f

OpenStudy (lgbasallote):

that would be 6! yes?

OpenStudy (lgbasallote):

since none repeated

OpenStudy (anonymous):

And how many 6 letter word from a,b,c,d,e,f,g

OpenStudy (zarkon):

now I'm getting sauravshakya answer...though I did it a different way

OpenStudy (anonymous):

Isnt it 7C6 *6!

OpenStudy (lgbasallote):

suppose so..yes

OpenStudy (anonymous):

Then what is wrong with my answer?

OpenStudy (lgbasallote):

it's not the same as your example

OpenStudy (anonymous):

I took four cases with no e , 1 e , 2 e and 3e

OpenStudy (lgbasallote):

what did you do for no e?

OpenStudy (anonymous):

7C6*6!

OpenStudy (lgbasallote):

for 1 e?

OpenStudy (anonymous):

7C5 *6!

OpenStudy (anonymous):

@lgbasallote ???

OpenStudy (zarkon):

36960 is correct (what sauravshakya got) though I wrote it as \[\sum_{x=0}^{3} {_7P_{6-x}}{6\choose x}\]

OpenStudy (anonymous):

It can also be 8C6 *6! + 8C5 *6!/2! + 8C4*6!/3!

OpenStudy (lgbasallote):

what about 2 e?

OpenStudy (lgbasallote):

@sauravshakya

OpenStudy (anonymous):

7C4 * 6!/2!

OpenStudy (lgbasallote):

ahh we had different answers in the 2 e

OpenStudy (lgbasallote):

i think you should divide that by 4...

OpenStudy (zarkon):

>> A=perms([1 2 3 4 5 6 7 8 8 8]); >> length(unique(A(:,1:6),'rows')) ans = 36960

OpenStudy (zarkon):

using octave

OpenStudy (lgbasallote):

wait...it seems i made a mistake in my solution...let me recheck

OpenStudy (lgbasallote):

yes. it's 46960

OpenStudy (lgbasallote):

36*

OpenStudy (zarkon):

octave would not lie ;)

OpenStudy (lgbasallote):

a simpler looking solution though is \[7! + 6C1 \times 7P5 + 6C2 \times 7P4 + 6C3 \times 7P2\]

OpenStudy (zarkon):

you fix your 7P2 (make it 7P3) then you have the solution I gave above

OpenStudy (lgbasallote):

yeah should be 7P3